OpenStudy (anonymous):

prove (tanx+sinx)/2tanx = cos^2(x/2)

OpenStudy (cwrw238):

try converting RHS to sin and cos

OpenStudy (anonymous):

I did, I'm still missing something.

OpenStudy (cwrw238):

i reduced it to (1 + cos x) --------- 2

OpenStudy (cwrw238):

i'll have to check std trig identities sorry i meant LHS in first post

OpenStudy (cwrw238):

i think that = RHS

OpenStudy (anonymous):

I have that also I have tried both sides, im missing a vital identity transformation

OpenStudy (anonymous):

So you were able to reduce LHS to (1+cos x)/2?

OpenStudy (cwrw238):

right cos^2 x = 1 + cos2x --------- 2

OpenStudy (cwrw238):

so cos^2 (x/2) = 1 + cos x -------- 2

OpenStudy (cwrw238):

its one of the half angle formulae

OpenStudy (anonymous):

LOL that was the easy part, Im trying to get (tanx+sinx)/2tanx to = that.

OpenStudy (cwrw238):

ok sin x + sin x ---- cos x -------------- 2 sin x ----- cos x = sinx + sinx cos x cosx --------------- * ----- cosx 2 sinx = sinx ( 1 + cos x) -------------- 2 sin x = ( 1 + cosx) ---------- 2

OpenStudy (anonymous):

Thank you very much, I did not factor correctly and ended up chasing my tail.

OpenStudy (cwrw238):

yw